# Handling Heterophily in Recommender Systems with Wavelet Hypergraph Diffusion

### v1

@@ 4 Evaluation
In this section, we assess the performance of WaveHDNN by comparing it with several state-of-the-art baseline methods. We begin by outlining the experimental setup and proceed to provide a detailed performance analysis across multiple datasets.
@@ 4.3 Performance Comparison
This section presents a detailed analysis of the performance results for all baseline models across the three datasets and provides a summary of their comparative performance.
@@
Model Amazon-books Recall@10 NDCG@10 Recall@20 NDCG@20 Recall@40 NDCG@40 LightGCN 0.05276 0.04472 0.08412 0.05566 0.13369 0.071 SGL 0.06331 0.05365 0.10035 0.06664 0.15863 0.08471 DHCF 0.06355 0.05345 0.09913 0.06576 0.15285 0.08254 HCCF 0.07332 0.06133 0.11695 0.07653 0.17942 0.09599 SHT 0.09123 0.07616 0.13929 0.09304 0.20774 0.11418 AutoCF 0.09765 0.0839 0.14391 0.09999 0.20815 0.12013 WaveHDNN 0.09837 0.08311 0.14714 0.09984 0.21345 0.12055 FWHDNN 0.10067 0.08517 0.15028 0.10254 0.21903 0.12394 %Improve. 2.33% 2.47% 2.13% 2.7% 2.61% 2.81%
@@
Model Steam Recall@10 NDCG@10 Recall@20 NDCG@20 Recall@40 NDCG@40 LightGCN 0.06503 0.06021 0.10555 0.07535 0.16418 0.09554 SGL 0.06541 0.06067 0.10524 0.07571 0.16385 0.09549 DHCF 0.07393 0.06773 0.11983 0.08515 0.18867 0.10842 HCCF 0.07756 0.0721 0.12503 0.08953 0.19201 0.11195 SHT 0.08006 0.07583 0.12687 0.02647 0.19513 0.11579 AutoCF 0.06562 0.07632 0.10627 0.07615 0.16475 0.09567 WaveHDNN 0.06547 0.07039 0.10071 0.08143 0.15383 0.09897 FWHDNN 0.06547 0.07039 0.10071 0.08143 0.15383 0.09897 %Improve. 4.36% 8.21% 13.9% 3.08% 18.72% 2.95%
@@
Model Yelp Recall@10 NDCG@10 Recall@20 NDCG@20 Recall@40 NDCG@40 LightGCN 0.05379 0.05072 0.08854 0.06302 0.14062 0.08065 SGL 0.0543 0.05103 0.08839 0.06308 0.14138 0.08093 DHCF 0.05302 0.04991 0.08855 0.06255 0.144 0.08154 HCCF 0.06265 0.05825 0.10426 0.07323 0.16882 0.09505 SHT 0.06955 0.06474 0.11281 0.08021 0.1836 0.10403 AutoCF 0.06694 0.06158 0.11401 0.07856 0.18773 0.10346 WaveHDNN 0.07393 0.06831 0.1212 0.08502 0.19427 0.10971 %Improve. 6.29% 5.51% 6.3% 5.99% 3.48% 5.45%
@@
Overall Comparison. As illustrated in Table 2 and Table 3, the FWHDNN model we propose consistently achieves superior performance compared to all baseline methods across various evaluation metrics, highlighting its robustness in diverse scenarios. Particularly, on the Steam dataset, which contains collection of user interaction data sourced from the Steam online gaming platform, our model exceeds the second-best model by a notable $2.95\%$ in NDCG@40, a critical metric for assessing the relevance of the top 40 recommendations. In addition, for sparse datasets, which feature limited user-item interactions, FWHDNN demonstrates marked improvements of $2.81\%$ and $5.45\%$, respectively, for Amazon-books and Yelp dataset over the next-best models. These gains are particularly noteworthy because sparse datasets typically pose greater challenges for recommendation models, as the lack of interactions makes it more difficult to learn accurate representations. Despite this, FWHDNN’s ability to generalize and perform well even in sparse environments underscores its adaptability and robustness.
@@
The superior performance of FWHDNN stems from several key innovations: 1) The ability to learn distinctive representations using hypergraph convolution networks with equivariant operators, effectively preserving both local and global neighborhood information. 2) The integration of multiple scales of localized features through a combination of hypergraph convolution and wavelet-based hypergraph transform layers, enabling more expressive feature extraction. 3) The adoption of intermediate and late fusion techniques to seamlessly combine structural and textual embeddings, enhancing the model’s capacity to capture diverse data modalities. These features collectively underscore FWHDNN’s adaptability to varying levels of interaction sparsity and dataset characteristics, establishing it as a powerful solution for collaborative filtering tasks. Overall, these advantages not only highlight the versatility of FWHDNN across diverse recommendation scenarios but also emphasize its capability to adapt to varying levels of interaction sparsity and dataset characteristics. The model’s multi-scale feature learning, coupled with its ability to preserve complex neighborhood relationships while considering multi-faceted features, makes it a powerful and effective tool for collaborative filtering tasks.
@@
Superiority over Hypergraph-based baselines. The experimental findings demonstrate that hypergraph-based approaches consistently surpass traditional graph-based models like LightGCN and SGL across all test scenarios. This performance difference underscores the benefits of leveraging hypergraphs to represent more intricate user-item relationships compared to standard binary interaction graphs. By capturing high-order group interactions that conventional graph methods often overlook, hypergraphs enable the generation of more precise recommendations.
@@
Notably, SHT, a representative hypergraph-based approach, outperforms SGL by $34.7\%$ and $21.2\%$ w.r.t. NDCG@40 in Amazon-books and Steam datasets, respectively. This significant improvement highlights the strength of hypergraphs in modeling complex group interactions and higher-order dependencies, leading to more accurate recommendations. By considering these high-order group interactions, hypergraph-based models can better represent the underlying topological structure of the user-item interaction space.
Among hypergraph-based models, FWHDNN excels with an $2.81\%$ gain in Recall@40 on the Amazon-books dataset, showcasing its exceptional capability in identifying relevant items for users. Recall@40, which measures a model’s ability to recommend items users are likely to interact with, reflects the effectiveness of FWHDNN’s sophisticated fusion of equivariant message passing and wavelet-based structural learning techniques. Such a substantial improvement in recall signifies that FWHDNN is highly effective at recommending items that closely match user preferences.
@@
In summary, the experimental results show the superiority of hypergraph-based methods over traditional graph-based approaches, especially when it comes to modeling complex, high-order relationships. Furthermore, among hypergraph-based models, our proposed WaveHDNN consistently achieves state-of-the-art performance, thanks to its advanced integration strategy of structural and textual information, multi-scale feature learning, and robust message-passing mechanisms.

### v2

@@ 4 Evaluation
In this section, we assess the performance of FWHDNN by comparing it with several state-of-the-art baseline methods. We begin by outlining the experimental setup and proceed to provide a detailed performance analysis across multiple datasets.
@@ 4.3 Performance Comparison
This section presents a detailed analysis of the performance results for all baselines across the three datasets.
@@
Model Amazon-books Recall@10 NDCG@10 Recall@20 NDCG@20 Recall@40 NDCG@40 LightGCN 0.05276 0.04472 0.08412 0.05566 0.13369 0.071 SGL 0.06331 0.05365 0.10035 0.06664 0.15863 0.08471 DHCF 0.06355 0.05345 0.09913 0.06576 0.15285 0.08254 HCCF 0.07332 0.06133 0.11695 0.07653 0.17942 0.09599 SHT 0.09123 0.07616 0.13929 0.09304 0.20774 0.11418 AutoCF 0.09765 0.0839 0.14391 0.09999 0.20815 0.12013 WaveHDNN 0.10132 0.08657 0.15136 0.10389 0.21812 0.12484 FWHDNN 0.10368 0.08871 0.15457 0.10670 0.22382 0.12834 %Improve. 2.33% 2.47% 2.13% 2.7% 2.61% 2.81%
@@
Model Steam Recall@10 NDCG@10 Recall@20 NDCG@20 Recall@40 NDCG@40 LightGCN 0.06503 0.06021 0.10555 0.07535 0.16418 0.09554 SGL 0.06541 0.06067 0.10524 0.07571 0.16385 0.09549 DHCF 0.07393 0.06773 0.11983 0.08515 0.18867 0.10842 HCCF 0.07756 0.0721 0.12503 0.08953 0.19201 0.11195 SHT 0.08006 0.07583 0.12687 0.02647 0.19513 0.11579 AutoCF 0.06562 0.07632 0.10627 0.07615 0.16475 0.09567 WaveHDNN 0.08607 0.07754 0.13525 0.09602 0.20651 0.12019 FWHDNN 0.08982 0.08391 0.15406 0.09900 0.24518 0.12374 %Improve. 4.36% 8.21% 13.9% 3.08% 18.72% 2.95%
@@
Model Yelp Recall@10 NDCG@10 Recall@20 NDCG@20 Recall@40 NDCG@40 LightGCN 0.05379 0.05072 0.08854 0.06302 0.14062 0.08065 SGL 0.0543 0.05103 0.08839 0.06308 0.14138 0.08093 DHCF 0.05302 0.04991 0.08855 0.06255 0.144 0.08154 HCCF 0.06265 0.05825 0.10426 0.07323 0.16882 0.09505 SHT 0.06955 0.06474 0.11281 0.08021 0.1836 0.10403 AutoCF 0.06694 0.06158 0.11401 0.07856 0.18773 0.10346 WaveHDNN 0.07393 0.06831 0.1212 0.08502 0.19427 0.10971 FWHDNN 0.07644 0.07238 0.13158 0.08750 0.21216 0.11286 %Improve. 3.39% 5.95% 8.56% 2.92% 9.21% 2.88%
@@
Overall Comparison. As illustrated in Table 2, Table 3, and Table 4, the proposed FWHDNN model consistently outperforms all baseline methods across the Amazon-books, Steam, and Yelp datasets, underscoring its robustness in diverse recommendation scenarios. On the Steam dataset, which comprises user interaction data from the Steam online gaming platform, FWHDNN achieves a notable improvement of $2.95\%$ in NDCG@40 compared to the second-best model, highlighting its effectiveness in capturing the relevance of the top-40 recommendations. On the Amazon-books dataset, which is relatively sparse, FWHDNN demonstrates consistent gains, including a $2.81\%$ improvement in NDCG@40. Similarly, on the Yelp dataset – another challenging sparse environment – FWHDNN yields substantial improvements, such as $8.56\%$ in Recall@20 and $9.21\%$ in Recall@40, outperforming all baselines. These results are particularly significant because sparse datasets with limited user-item interactions often hinder representation learning, yet FWHDNN adapts well and delivers consistently strong performance across all benchmarks.
@@
The superior performance of FWHDNN is driven by three key innovations: (1) its ability to learn distinctive representations through hypergraph convolution networks with equivariant operators, preserving both local and global neighborhood information; (2) the integration of multiple scales of localized features via a combination of hypergraph convolution and wavelet-based hypergraph transform layers, enabling richer and more expressive feature extraction; and (3) the adoption of intermediate and late fusion strategies to combine structural and textual embeddings, thereby capturing diverse data modalities.
@@
Superiority over Hypergraph-based baselines. The experimental findings demonstrate that hypergraph-based approaches consistently surpass traditional graph-based models such as LightGCN and SGL across all datasets. This consistent performance gap underscores the benefits of using hypergraphs to represent more intricate user–item relationships compared to standard binary interaction graphs. By capturing high-order group interactions that conventional graph methods overlook, hypergraph-based models can generate more precise and robust recommendations.
@@
Notably, SHT, a representative hypergraph-based approach, outperforms SGL by $34.8\%$, $21.2\%$, and $29.6\%$ with respect to NDCG@40 on the Amazon-books, Steam, and Yelp datasets, respectively. These substantial improvements highlight the strength of hypergraphs in modeling complex group interactions and higher-order dependencies, leading to more accurate and reliable recommendation outcomes.
Among hypergraph-based models, FWHDNN stands out by further pushing the performance boundaries. On the Amazon-books dataset, FWHDNN achieves a $2.81\%$ gain in NDCG@40 over the next-best baseline. On the Steam dataset, FWHDNN delivers a $2.95\%$ improvement in NDCG@40, while on the Yelp dataset, it achieves notable gains of $8.56\%$ and $9.21\%$ in Recall@20 and Recall@40, respectively. These consistent improvements across different domains – including both dense (Steam) and sparse (Amazon-books, Yelp) datasets – demonstrate FWHDNN’s adaptability and effectiveness in capturing user preferences under varying levels of interaction sparsity.
@@
In summary, the results confirm the superiority of hypergraph methods over traditional graph-based approaches, especially for modeling complex, high-order dependencies in recommendation tasks. Moreover, among hypergraph-based approaches, FWHDNN consistently achieves state-of-the-art performance, owing to its advanced integration of structural and textual information, multi-scale feature learning, and robust message-passing mechanisms.
Table 5: Ablation test of FWHDNN (in bold) with its variants.
Ablation Settings Amazon-books Steam Yelp Recall@40 NDCG@40 Recall@40 NDCG@40 Recall@40 NDCG@40 FWHDNN 0.22382 0.12834 0.24518 0.12374 0.21216 0.11286 w/o Cross-Difference Relation Encoder 0.18623 0.10627 0.19123 0.1104 0.17236 0.09499 w/o Multi-level Cluster-wise Encoding 0.21035 0.11626 0.18845 0.10701 0.16856 0.09368 w/o Integrated Multi-modal Fusion Mechanism 0.21459 0.10489 0.21024 0.11014 0.17565 0.08448
4.4 Ablation Test
In this section, we assess the significance of the key components of our FWHDNN. We create variation models by replacing or removing each component, demonstrating the performance improvements achieved by the proposed elements. The statistics of results are summarized Table 5.
Cross-Difference Relation Encoder. This encoder enhances the model’s ability to capture diverse user–item relationships by incorporating heterophily into the graph structure. As shown in Table 5, removing this component leads to a clear drop in performance across all datasets (e.g., Amazon-books: Recall@40 falls from 0.22382 to 0.18623). The encoder’s adaptive message-passing mechanism within HGCN layers helps preserve diversity and mitigates over-smoothing, resulting in more robust embeddings and improved recommendation accuracy.
Multi-level Cluster-wise Encoding. The wavelet-based HGCN layer captures multi-scale graph structures, ensuring that both global and fine-grained neighborhood patterns are preserved. Eliminating this mechanism causes a noticeable degradation, as reflected in the Steam dataset where Recall@40 decreases from 0.24518 to 0.18845. These results emphasize its critical role in maintaining structural detail across scales. By aggregating information at multiple levels, the model avoids over-smoothing and achieves more accurate collaborative filtering.
Integrated Multi-modal Fusion Mechanism. This mechanism combines heterogeneous feature spaces, enabling the model to leverage complementary modalities for richer representation learning. As Table 5 shows, excluding this component produces consistent performance losses (e.g., Yelp: NDCG@40 drops from 0.11286 to 0.08448). The results highlight the importance of effective fusion in capturing inter-modal signals, ensuring that FWHDNN can fully exploit structural and contextual information. This synergy across modalities underpins the model’s superior ability to generalize and deliver robust predictions.
4.5 Hyper-parameter Studies
This experiment examines how sensitive our model is to different hyperparameters. Among the commonly used hyperparameters, we specifically focus on the number of HGCN layers and the embedding dimension, as these factors have a notable influence on performance outcomes.
Figure 3: Impact of the number of HGCN layers and different embedding size.
Impact of HGCN Layer Depth. We evaluated the effect of varying the number of HGCN layers on FWHDNN’s performance, as shown in Figure 3 (left). The number of layers was varied from 1 to 5 across the Amazon-books, Steam, and Yelp datasets. The results reveal a consistent trend: performance improves as the number of layers increases up to three, after which it slightly declines. For instance, Amazon-books achieves its highest NDCG@40 at three layers, while both Steam and Yelp show similar peak behavior before a mild reduction at deeper depths.
This pattern suggests that moderate depth allows the model to capture meaningful higher-order relationships without oversmoothing. However, adding more layers introduces noise from overly distant neighbors, which can dilute useful signals. These findings highlight the importance of choosing an appropriate number of layers – in this case, three – as a balance between local connectivity and global structural information.
Impact of Embedding Dimension. We further examined the influence of embedding size, varying the dimensionality across [8, 16, 32, 64, 128]. As depicted in Figure 3 (right), performance improves as the dimension increases to 32, where the highest NDCG@40 values are observed across all three datasets. Beyond this point, accuracy remains stable or slightly decreases. For example, Amazon-books peaks at 32 dimensions, while Steam and Yelp exhibit similar trends with only marginal gains at larger sizes.
The diminishing returns at higher dimensions suggest that excessively large embeddings may spread information too thinly across dimensions, reducing the model’s ability to focus on meaningful patterns. Moreover, larger embeddings impose additional computational and memory overheads without commensurate improvements in accuracy. Thus, selecting a moderate embedding size, such as 32, is optimal to ensure both efficiency and robust performance in recommendation tasks.